OpenStudy (anonymous):

Solve (8y – 7) – (5y + 1) = 4. y = 10/3 y = -2/13 y = 16/13 y = 4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

remove parentheses. combine like terms add or subtract to get y on one side by itself. divide to get y by itself

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

3y=12 y=4

OpenStudy (anonymous):

\[(8y-7)-(5y+1)=4\] \[8y-7-5y-1=4\] \[3y-8=4\] \[3y=12\] \[y=\frac{12}{3}\] \[y=4\]

OpenStudy (anonymous):

(8y-5y)-(7+1)=4 3y-8=4 3y=12 y=4

OpenStudy (anonymous):

(8y-7) - (5y+1) = 4 8y-7 - 5y-1 = 4 3y - 8 = 4 3y = 4+8 3y = 12 y =12/3 y = 4

OpenStudy (anonymous):

understood